OpenStudy (anonymous):

find the perimeter of a lot with sides measuring 2sqrt8, 5sqrt18, 3sqrt50, and 5sqrt75

OpenStudy (ganpat):

whats perimeter over here ??

OpenStudy (anonymous):

Huh?

OpenStudy (ganpat):

perimeter = sum of all sides... write this in equation form.. meaning, add all the sides... (we will go together..)

OpenStudy (ganpat):

4 sq rt 2 + 15 sq rt 2 + 15 sq rt 2 + 25 sq rt 3

OpenStudy (anonymous):

where did the 15sqrt2 come from?

OpenStudy (ganpat):

34 sq rt 2 + 25 sq rt 3..

OpenStudy (ganpat):

18 = 9 * 2 = 3 sq * 2 50 = 25 * 2 = 5 sq * 2

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

34sqrt2 + 25sqrt3